lolokrat:

can someone explain what is going on in this problem? what exactly am i solving for? how do i know that x does not equal 1 or -1?

lolokrat:

lolokrat:

none of this was lectured in my class or presented in my homework, so im pretty lost here. what can i put into youtube to find out how to work these steps LOL

mxddi3:

Wait you're on number 4 right?

lolokrat:

yes i believe it's 2 parts

mxddi3:

Well to start, we have \[\frac{ x+3 }{ (x-1)(x+1) }\] for problems like this, our denominator can NOT BE EQUAL TO ZERO. So, we must make the denominator into its own problem by setting it equal to 0. \[(x-1)(x+1)=0\] now, we can make both sets of parentheses equal to 0, meaning we end up with: (x-1)=0 and (x+1)=0 Now, we solve to get x by itself. For the first one we add 1 to both sides giving us x=1. For the second we will subtract x from both sides giving us x=-1. Now remember, the denominator CAN NOT be equal to 0. In our problem, we set the denominator equal to 0, meaning we just solved to find the values that x CAN NOT BE. Therefore, \[x \neq1 \] and \[x \neq-1\] does this make sense?

lolokrat:

oh my god what lol you did an excellent job of describing it but my brain ?????? what is this type of problem? why?? what exactly am i solving for??

lolokrat:

and what purpose does the numerator x+3 serve in the problem?

mxddi3:

okay so you are looking for what x can not be equal to. I learned something very similar to this a few weeks back. I think we're looking for the domain of the function. in this problem, we are not doing anything with the numerator of x+3. We are simply looking for what x could not be for the domain i'm pretty sure. When you see these problems in this lesson, we leave the numerator be and we simply just have to make sure our denominator does not end up being equal to zero. let me see if i could find you a video to explain. Khan academy does work very well for teaching these.

lolokrat:

ah ok that makes a bit more sense

mxddi3:

https://www.youtube.com/watch?v=djT6-YamHaA try watching the first i'd say 5 minutes of this. It should do a good job showing you just in case you're still not understanding.

lolokrat:

thank you, the video did help a bunch. so how did they come up with the HA y=o?